# The Combinatorics of Motzkin Polyominoes

### v1

@@ 6. The interior points statistic
In this section, we study the statistic of the number of interior points on Motzkin polyominoes. As for the previous section, for all $1\leq i\leq n$, ${\mathcal{M}}_{n,i}$ is the set of the Motzkin words of length $n$ whose last symbol is $i$, and we define the generating functions
@@
By multiplying by $v^{i-1}$ the last equation and summing over $i\geq 2$, we obtain the functional equation
@@
By differentiating twice this equation with respect to $v$ and taking $v=xM(x)+1$, we obtain that
@@
As an application of Pick’s theorem (cf. [6, pp. 40]), one may establish a relation between the area, semiperimeter, and number of interior points of a Motzkin polyomino. Pick’s theorem says that the area of a simple polygon with integer vertex coordinates is equal to $I+B/2-1$, where $I$ is the number of interior points and $B$ is the number of points lying on the boundary. Let $\texttt{int}(n)$ be the sum of the interior points over all Motzkin polyominoes of length $n$.
@@
Acknowledgments. José L. Ramírez was partially supported by Universidad Nacional de Colombia. The third author would like to thank the Laboratoire d’Informatique de Bourgogne for the warm hospitality during his visit at Université de Bourgogne where part of this work was done. This work was supported in part by ANR-22-CE48-0002 funded by l’Agence Nationale de la Recherche and by the project ANER ARTICO funded by Bourgogne-Franche-Comté region (France).

### v3

@@ 6. The interior points statistic
In this section, we study the statistic of the number of interior points on Motzkin polyominoes. As in the previous section, for all $1\leq i\leq n$, ${\mathcal{M}}_{n,i}$ is the set of the Motzkin words of length $n$ whose last symbol is $i$, and we define the generating functions
@@
By multiplying the last equation by $v^{i-1}$ and summing over $i\geq 2$, we obtain the functional equation
@@
By twice differentiating this equation with respect to $v$ and taking $v=xM(x)+1$, we obtain that
@@
As an application of Pick’s theorem (cf. [7, pp. 40]), one may establish a relation between the area, semiperimeter, and number of interior points of a Motzkin polyomino. Pick’s theorem says that the area of a simple polygon with integer vertex coordinates is equal to $I+B/2-1$, where $I$ is the number of interior points and $B$ is the number of points lying on the boundary. Let $\texttt{int}(n)$ be the sum of the interior points over all Motzkin polyominoes of length $n$.
@@
Acknowledgments. The authors are grateful to the referees for the detailed comments and corrections that helped us improve the paper. José L. Ramírez was partially supported by Universidad Nacional de Colombia. The third author would like to thank the Laboratoire d’Informatique de Bourgogne for the warm hospitality during his visit at Université de Bourgogne where part of this work was done. This work was supported in part by ANR-22-CE48-0002 funded by l’Agence Nationale de la Recherche and by the project ANER ARTICO funded by Bourgogne-Franche-Comté region (France).